# The Hands-Up Problem and How to Deal With It: Secondary School Teachers’ Experiences of Debugging in the Classroom

### v1

@@ 4.5 Improving students’ debugging ability
As well as general scaffolding, the majority of teachers reported using teaching approaches that are designed to help students become more successful debuggers. A more general sub-theme of this theme is teaching programming strategies to make debugging easier. These are patterns of programming that prevent many errors from accumulating, or help keep a student aware of the purpose of their program. Adam and Hannah, for example, encouraged an iterative process of small changes and testing.
@@
A few teachers were strong advocates of utilising programming error messages (PEMs). Although they are not always intuitive to understand, Adam and Felix viewed PEMs as valuable sources of immediate feedback.
@@
This positive outlook on PEMs drives these teachers to frequently discuss them. The act of ‘decoding’ PEMs was subsequently mentioned by some of these teachers.
Through repeated exposure and practice, guided by the strategies they teach, teachers do observe positive change in their students’ debugging behaviours. Benjamin’s students, who are ‘pre-cued to look for common errors’, become drilled in doing this themselves.
@@
This positive change appears to be more frequently reported by the more confident teachers we interviewed. Regardless, such improvement is to be celebrated. These teachers show how their teaching strategies help students to go about it with independence and confidence. With this comes more of the aforementioned elation of solving an error, as well as positive initial experiences for students learning program.
@@ 4.6 Promoting a positive error culture
Students’ improvement in debugging is not solely due to instructional strategies. Cultivating a positive error culture, an environment where students feel comfortable making errors in their programs, also helps students become more resilient to errors. One useful method for promoting a positive error culture is getting used to erroneous code. This may involve sabotaging activities (Benjamin and Hannah), where students actively add errors to their peers’ programs, or giving students erroneous code (Chloe and Gideon). Hannah points out that sabotaging is useful because it exposes students to errors in a depersonalised format, reducing the potential emotional angst around debugging.
“It doesn’t matter if it doesn’t work. It’s not yours. And so that was a really nice way to kind of break down the anxiety about that there will be errors.” - Hannah
Another method of normalising errors in the classroom is by highlighting that errors are to be expected rather than dreaded. Felix explains how ‘failures’ should not just be reported, but also celebrated to remove the stigma around them.
@@
Felix’s response also highlights the practice of celebrating the effort as well as the outcome. If students are actively rewarded for sustained debugging effort, as also in Adam’s case, this can act as a further motivator for teacher-independent debugging.
@@
These approaches, combined with verbal encouragement throughout the debugging process, help create an environment where both encountering and struggling with errors are natural experiences. Such measures help make debugging a healthy, normal activity that can be a source of joy.
@@ 5 Discussion
Teachers’ experiences of debugging in the classroom spanned a range of insights and issues. Many of these align with the related work discussed in Section 2, particularly that related to the teacher-student interaction (Section 2.2.1). While these studies focus on how students assist a particular student(s), the teachers we interviewed shared more general experiences and feelings with debugging in the classroom.
A major motif that we found in teachers’ experiences is the common reliance students have on their teacher for debugging support. This tended to be the case for less motivated students, who would raise their hands at the point of or soon after encountering an error. We think much of our results, and some other reported teachers’ experiences, can be conceptualised through a phenomenon we call the “the hands-up problem”.
@@ 5.1 The Hands-Up Problem
The hands-up problem is the in-class difficulty deriving from the common student tendency to raise their hand when struggling to complete a task. Since the need for teacher support is reportedly common (see the commonness of teacher reliance and Bagge (2015); Michaeli and Romeike (2019a); Uppal (2020); Kerslake (2024)), the problem derives from many students raising their hands at once. Therefore, the severity of the problem scales with the size of teachers’ classrooms, making it more difficult for typical school teachers whose classes contain up to 30 students (Yadav et al., 2016a).
This phenomenon may indeed happen in any subject; hand-raising is a common classroom behaviour. However, many studies outside of computing education investigate hand-raising and student engagement in the context of whole class discussions (e.g., (Turner and Patrick, 2004; Böheim et al., 2020a, b)). The dynamics of motivation, emotion, and classroom atmosphere in this setting are different to when students are programming themselves. We therefore focus on the context of debugging.
According to the teachers we interviewed, students generally raised their hands at the point of encountering an error. However, hand-raising may occur after students have put some initial effort into debugging (Michaeli and Romeike, 2019a; Wachter and Michaeli, 2025). Either way, the consequences for teachers and students are generally negative. We now describe some drivers, consequences, and counters to the hands-up problem, visualised in Figure 1.
@@ 5.1.1 The Drivers
There are a range of factors that prompt a student to raise their hands during debugging struggle, which we call drivers. These drivers may be domain-agnostic or domain-specific. This list is not meant to be exhaustive, but aims to give a general categorisation of some factors at play.
Domain-agnostic drivers for hand-raising are factors that can contribute to the hands-up problem in any subject. We believe that these are generally related to students’ affect. Teachers in our study often discussed a general lack of resilience to work through a problem, leading to reliance on the teacher. This problem is also highlighted by teachers in Sentance and Csizmadia (2017); Wachter and Michaeli (2025) and likely relates to students’ motivation levels. However, the very act of hand-raising must indicate some motivation to progress, otherwise students would disengage (see negative emotion during debugging struggle). Therefore, drivers at play here may be the motivation to problem-solve independently of the teacher, or attitudes of low self-efficacy or self-reliance.
Several domain-specific drivers for debugging may also contribute to the hands-up problem. These include difficulties specific to Python and learning to program more widely, such as the difficulty of getting syntax exactly right and conceptual difficulties. Lots of research has already reported on common beginner programmers’ errors and misconceptions (e.g., (Du Boulay, 1986; Pea, 1986; Sorva, 2021)), including the ‘syntax barrier’ beginners face (Denny et al., 2011, 2012). The size of these research areas indicates the number of potential stumbling blocks that may contribute to a student seeking teacher support. Difficult to interpret PEMs are also certainly a driver, which relates to the active field of PEM readability with computing education research (e.g., (Becker et al., 2019; Denny et al., 2021; Phung et al., 2023)).
Generally, domain-specific drivers can be viewed through cognitive load theory (Sweller, 1988). If a student’s working memory is near-saturated when programming, errors and difficulty with problem-solving follow. If this is the case, hand-raising is arguably a logical behaviour, because they lack the cognitive capacity to debug themselves. However, this has the side-effect of contributing towards the hands-up problem if the intrinsic cognitive load of the task in question is too high.
@@ 5.1.2 The Consequences
Most teachers in our study alluded to many students raising their hands for debugging support, particularly when starting to learn to program. This is why we call this hands-up phenomenon a problem; in typical classroom settings, teachers simply cannot go around and effectively support every student (Yadav et al., 2016a; Michaeli and Romeike, 2019a). This has both student-side and teacher-side consequences.
The student-side consequences are multi-faceted. Once students raise their hands, teachers in our study described how unlikely it was for students to progress until they come to help. Given the drivers discussed, this is not surprising, as students may not have the working memory to continue their debugging efforts. As a result, some students will linger in a state of being stuck, or emotional angst, for an extended period. Even when a teacher does come to provide support, the nature of the interaction will be altered if there is a time pressure (Elliott et al., 2024). Such a classroom atmosphere makes it very difficult to help students learn from an error they are struggling with.
The emotional consequences of debugging are also important to consider. Feelings of frustration, anxiety, and disappointment are documented in this study (the emotional nature of learning to debug) and previous work (Kinnunen and Simon, 2010; Gorson et al., 2022; Gale and Sentance, 2023; Ryan and DeLiema, 2023). If these feelings occur repeatedly, students may simply disengage in future encounters with errors (see negative emotion during debugging struggle). If these experiences are students’ first with learning to program, such emotional arousal and lack of success can negatively form students’ programming self-efficacy (Bandura, 1978; Gorson and O’Rourke, 2020). Such experiences are therefore important to limit, but hard to do so if students are reliant on teacher support.
There are also teacher-side consequences of the hands-up problem. Some teachers described their classroom atmosphere as ‘chaotic’ and ‘busy’ when lots of hands are raised. Some also reported feeling a duty to help every student requesting assistance. Failing to do so can evoke feelings of inadequacy or disappointment (see the toll on the teacher of this reliance). Furthermore, if teachers are regularly assessed or inspected, a sea of raised hands may give the impression of an unproductive lesson. This again can evoke negative feelings in teachers, who may over-compensate by providing too much assistance to students. In Uppal (2020)’s reflections on their in-class debugging practice, “in the midst of a busy classroom where as a teacher I was multitasking on many levels, it was all too easy to over-scaffold and provide too much direction, to give the appearance of a seemingly successful lesson” (p.15).
@@ 5.1.3 The Counters
So far, we have characterised the hands-up problem in fairly negative terms. However, it can be counteracted. Several teachers in our study (see motivating teacher-independent debugging) and related work (Bagge, 2015; Michaeli and Romeike, 2019a; Wachter and Michaeli, 2025) report using teaching approaches that encourage teacher-independent debugging. Like the drivers, these counters can be divided into domain-agnostic techniques and domain-specific counters.
@@
Domain-agnostic counters are strategies that teachers of any subject can apply to alleviate the hands-up problem. For instance, teachers may refuse to help students until they evidence some attempt at teacher-independent problem-solving (see refusing to help students and Bagge (2015); Michaeli and Romeike (2019a)). Some may also encourage collaboration with peers before providing support (see encouraging helpful collaboration).
@@
Debugging-specific counters include referring students to resources and tools that aid successful debugging (see signposting to resources for assistance and Nixon et al. (2023); Ryan and DeLiema (2023); Elliott et al. (2024)) and modelling the debugging process to the whole class (see Guiding students through debugging strategies and Ryan and DeLiema (2023)). An obvious benefit of these approaches is that a whole class can benefit from them simultaneously. Teachers may also teach heuristics for ‘decoding’ PEMs due to their oft-ambiguity (utilising programming error messages). An attitudinal counter also mentioned by some teachers is the enforcement of a positive error culture (Hennessy Elliott et al., 2023). When students learn to program in an environment where errors are considered normal, they will likely feel more confident to attempt independent debugging.
@@
These strategies may take time for students to become accustomed to (Bagge, 2015). However, the teachers who report consistently using these counters also report their students becoming more perseverant and less reliant on the teacher when debugging (Michaeli and Romeike, 2019a). Ultimately, this gives the teacher more time to provide more meaningful support to a smaller number of struggling students.
@@
From our sample of teachers, teachers who reported these counters tended to be more confident and have decades of experience. This is less the case for the less confident teachers, leaving them more susceptible to the consequences of the hands-up problem. More generally, how widespread such counters are in computing classrooms is unclear. Results from Sentance and Csizmadia (2017) suggest only a fraction of teachers have strategies for teaching debugging and resilience to errors, with lack of time to teach programming probably a limiting factor.
@@ 5.2 The Implications
Given the increasing volume of computing education research focused on debugging, it is surprising that some teachers report struggling with the hands-up problem. Although we only interviewed nine teachers, we expect our findings align with more teachers’ experiences due to the commonality of non-specialist computing teachers and reported lack of content knowledge by some (Yadav et al., 2016a; Sentance and Csizmadia, 2017). For teachers who report using counters, experience or professional development seemed to be the reason. This begs the question: how can less experienced teachers [^7] This includes both newly qualified and non-specialist computing teachers. more confidently teach debugging? Or in terms of Figure 1, how can counters to the hands-up problem be developed? We did not identify any themes related to these questions in our analysis and therefore only provide an initial recommendation for future research.
An obvious suggestion is to improve less experienced teachers’ debugging PCK. Despite its integrity to programming and influence on classroom practice, we are not aware of a conceptualisation of debugging PCK in the research literature. Admittedly, a debugging PCK could already exist in debugging teaching resources or PD, but investigating this is beyond the scope of this study.
A debugging PCK would inevitably have some commonalities with conceptualisations of programming PCK (Saeli et al., 2011), such as the motivation for teaching it and common student difficulties. But it would additionally include debugging-specific pedagogy, techniques for effectively managing teacher-student interactions, and other aforementioned counters to the hands-up problem. Therefore, even if a teacher does not have time to explicitly teach the debugging process, a more developed debugging PCK would still equip them to deal with the hands-up problem.
@@
Once a concept of debugging PCK is more well-developed, it can be spread through professional development. Some studies have already reported successful debugging-specific PD (Elliott et al., 2022; Michaeli and Romeike, 2022; Nixon et al., 2023), which provides a good starting point for further development. However, given some teachers still report findings aligned with the hands-up problem (Bagge, 2015; Sentance and Csizmadia, 2017; Michaeli and Romeike, 2019a; Uppal, 2020; Kerslake, 2024), debugging-specific PD is either not widespread enough or not being effectively translated into practice. If the former is the case, more debugging-related PD needs to be developed or expanded. This does not need to solely focus on debugging; more emphasis on debugging could be included in programming PD workshops, given the intertwined nature of these skills.

### v3

@@ 4.5 Improving students’ debugging ability
As well as general scaffolding, most teachers reported using approaches designed to help students become more successful debuggers. A more general sub-theme of this theme is teaching programming strategies to make debugging easier. These are patterns of programming that prevent errors from accumulating or help to keep a student aware of the purpose of their program. Adam and Hannah, for example, encouraged an iterative process of small changes and testing.
@@
A few teachers were strong advocates of utilising programming error messages (PEMs). Although they are not always intuitively phrased, Adam and Felix viewed PEMs as valuable sources of immediate feedback.
@@
This positive outlook on PEMs drives these teachers to frequently discuss them. Some teachers also mentioned teaching students how to ‘decode’ PEMs.
Through repeated exposure and practice, guided by the strategies they teach, teachers reported a positive change in their students’ debugging behaviours. Benjamin’s students, who are “pre-cued to look for common errors”, become drilled in doing this themselves.
@@
This positive change appears to be more frequently reported by the more confident or experienced teachers we interviewed. Regardless, such improvement is to be celebrated. These teachers show how their teaching strategies help students to go about debugging with independence and confidence. With this comes more of the aforementioned elation of solving an error, as well as positive initial experiences for students learning to program.
@@ 4.6 Promoting a positive error culture
Students’ improvement in debugging is not solely due to instructional strategies. Cultivating a positive error culture, an environment where students feel comfortable making errors in their programs, also helps students become more resilient when debugging. One useful method for promoting a positive error culture is getting used to erroneous code. This may involve sabotaging activities (Benjamin and Hannah), where students actively add errors to their peers’ programs, or giving students erroneous code (Chloe and Gideon). Hannah points out that sabotaging is useful because it exposes students to errors in a depersonalised format, reducing the potential emotional angst around debugging.
“‘It doesn’t matter if it doesn’t work. It’s not yours. And so that was a really nice way to kind of break down the anxiety about that ‘there will be errors’.” - Hannah
Another method of normalising errors in the classroom is by highlighting that errors are to be expected rather than dreaded. Felix explains how ‘failures’ should not just be reported, but celebrated to remove the stigma around them.
@@
Felix’s response also highlights the practice of celebrating the effort as well as the outcome. If students are actively rewarded for their sustained debugging efforts, as in Adam’s case, this can act as a further motivator for teacher-independent debugging.
@@
These approaches, combined with verbal encouragement throughout the debugging process, help to create an environment where both encountering and struggling with errors are natural experiences. Such measures help make debugging a healthy, normal activity that can be a source of joy rather than despair.
@@ 5 Discussion
Teachers’ experiences of debugging in the classroom spanned a range of insights and issues. A major motif we found among these experiences was the common reliance students have on their teacher for debugging support. This tended to be the case for less motivated students, who would raise their hands at the point of or soon after encountering an error.
Our findings, along with previous reports of teachers’ experiences with debugging in the classroom, led us to conceptualise the hands-up problem, the in-class difficulty deriving from the common student tendency to raise their hand when struggling to complete a task. Although our definition frames the hands-up problem as a potential challenge for teachers of any subject, we focus on programming classrooms due to the lack of similar studies with other subject teachers. Therefore, in this paper, ‘struggling to complete a task’ refers to struggling to debug when learning to program.
We now introduce the hands-up problem as a challenge that computing teachers face in their programming lessons as a result of students struggling with debugging. We also propose ways to alleviate the problem. In addition to prior work, we reference theme names from the results section in italics where appropriate.
@@ 5.1 The Hands-Up Problem
The hands-up problem is the in-class difficulty deriving from the common student tendency to raise their hand when struggling to complete a task. Since the need for teacher support is reportedly common (see the commonness of teacher reliance and CombattingLearnedHelplessness; CurrentPerspectivesDebugging; LackPerserveranceAutonomyDebugging; StumpTheTeacher), the problem derives from many students raising their hands at once. Therefore, the severity of the problem scales with the size of the classroom, making it difficult for school teachers whose classes contain 25-30 students (ExpandingCSChallengesExperiences; UKAverageClassSizes).
According to the teachers we interviewed, students generally raised their hands at the point of encountering an error. Hand-raising may also occur after students have put some initial effort into debugging (CurrentPerspectivesDebugging; TeacherDiagnosticInterventionSkills). Either way, the consequences for teachers and students are generally negative. We now describe some drivers, consequences, and counters to the hands-up problem, visualised in Figure 1.
@@ 5.1.1 The Drivers
There is a range of drivers that prompt a student to raise their hands during debugging struggle. These may be domain-agnostic or domain-specific. The list below is not meant to be exhaustive, but gives a general categorisation of some factors at play.
Domain-agnostic drivers for hand-raising are factors that can contribute to the hands-up problem in any subject. We argue that these are generally related to students’ affect. Teachers in our study often discussed a general lack of resilience to work through a problem, leading to reliance on the teacher. This problem is also highlighted by teachers in related work (ComputingInCurriculumChallengesStrategies; TeacherDiagnosticInterventionSkills) and likely relates to students’ motivation levels. However, the very act of hand-raising must indicate some motivation to progress, otherwise students would disengage (see negative emotion during debugging struggle). Therefore, the drivers at play here may be a lack of motivation to problem-solve independently of the teacher, or attitudes of low self-efficacy.
Several domain-specific drivers for debugging may also contribute to the hands-up problem. For example, students learning a block- or text-based programming language may have misconceptions (see conceptual difficulties and (MisconceptionsBeginnerProgrammer)) that make errors difficult to spot without the help of a more knowledgeable other. Teachers in our study also discussed drivers specific to learning a TBPL, such as the difficulty of getting syntax exactly right and difficult to interpret PEMs. Again, previous work has repeatedly identified these as difficulties with debugging in a TBPL (UnderstandingSyntaxBarrier; AllSyntaxErrorsNotEqual; ErrorMessageWorkingGroup). We suggest how these could be alleviated in Section 5.1.3.
Generally, domain-specific drivers can be viewed through cognitive load theory (OriginalCognitiveLoadArticle; CognitiveArchitecture20YearsLater). If a student’s working memory is near-saturated when programming, errors and difficulty with problem-solving follow (OriginalCognitiveLoadArticle). Hand-raising is thus a logical behaviour, as students lack the cognitive capacity to debug themselves. However, if many students are experiencing a high load on working memory, the hands-up problem quickly follows.
@@ 5.1.2 The Consequences
Most teachers in our study alluded to many students raising their hands for debugging support, particularly when first learning to program. This is why we call it the hands-up problem; in typical classroom settings, teachers simply cannot go around and effectively support every student (ExpandingCSChallengesExperiences; CurrentPerspectivesDebugging). This has both student-side and teacher-side consequences.
The student-side consequences are multi-faceted. Once students raise their hands, teachers in our study explained how unlikely it was for students to progress until they came to help. Given the drivers discussed, this is not surprising, as students may not have the working memory to continue their debugging efforts. However, this means some students will linger in a state of being stuck or emotional angst for an extended period. Even when a teacher does come to provide support, the teacher-student interaction will likely be rushed if there is a time pressure or other students with their hands up (CurrentPerspectivesDebugging; TeacherSupportForDebuggingPhysicalComputing). Such a classroom atmosphere makes it very difficult for teachers to meaningfully guide students through the errors they are struggling with.
The emotional consequences of debugging are also important to consider. Feelings of frustration, anxiety, and disappointment are documented in this study (the emotional nature of learning to debug) and previous work (ProgrammingAssignmentsEmotionalToll; EDAProgrammingEmotions; ReflectionsSustainedDebuggingSupport). If these feelings occur repeatedly, students may simply disengage in future encounters with errors (see negative emotion during debugging struggle). If experiences of emotional arousal and lack of success are students’ first with learning to program, they can negatively form students’ programming self-efficacy (OriginalSelfEfficacyArticle; ProgrammingSelfEfficacySelfAssessment). Such experiences are therefore important to limit, but hard to do so if students are heavily reliant on teacher support.
There are also teacher-side consequences of the hands-up problem. Some teachers described their programming lessons as ‘chaotic’ and ‘busy’, in part due to the number of hands raised. Some also reported feeling a duty to help every student requesting assistance. Failing to do so evoked feelings of inadequacy or disappointment (see the toll on the teacher of this reliance). Furthermore, if teachers are regularly assessed or inspected, a sea of raised hands may give the impression of an unproductive lesson. This can again evoke negative feelings in teachers, who may overcompensate by providing too much assistance to students. As LackPerserveranceAutonomyDebugging states in their reflections on their in-class debugging practice, “in the midst of a busy classroom where as a teacher I was multitasking on many levels, it was all too easy to over-scaffold and provide too much direction, to give the appearance of a seemingly successful lesson” (LackPerserveranceAutonomyDebugging, p. 15).
@@ 5.1.3 The Counters
So far, we have characterised the hands-up problem in fairly negative terms. However, it can be counteracted. Several teachers in our study report using teaching approaches that encourage teacher-independent debugging (see motivating teacher-independent debugging and CombattingLearnedHelplessness; CurrentPerspectivesDebugging; TeacherDiagnosticInterventionSkills). We refer to these as counters to the hands-up problem. Like the drivers, the counters can be divided into domain-agnostic and domain-specific.
@@
Domain-agnostic counters are strategies that teachers of any subject can apply to alleviate the hands-up problem. For instance, teachers may refuse to help students until they evidence some teacher-independent problem-solving (see refusing to help students and CombattingLearnedHelplessness; CurrentPerspectivesDebugging). Some may also encourage collaboration with peers before providing support (see encouraging helpful collaboration).
@@
There are also a variety of domain-specific counters that support students with the debugging process. Teachers reported referring students to appropriate resources and tools (see signposting to resources for assistance and SupportingStudentsScienceInquiry; ReflectionsSustainedDebuggingSupport; TeacherSupportForDebuggingPhysicalComputing) and modelling the debugging process to the whole class (see guiding students through debugging strategies and ReflectionsSustainedDebuggingSupport). Some also taught heuristics for ‘decoding’ PEMs due to their often-ambiguous nature (see utilising programming error messages). An obvious benefit of these counters is that the whole class can benefit from them simultaneously. An attitudinal counter mentioned by some teachers is the enforcement of a positive error culture (see promoting a positive error culture and TowardDebuggingPedagogy).
As well as the domain-specific counters that teachers mentioned, debugging-specific teaching approaches developed in prior work could be useful counters. Systematic debugging processes (e.g., (CarverImprovingChildrensDebugging; SystematicProcessMichaeli)) or debugging-specific tooling (e.g., (PythonTutor; Ladebug; SystematicDebuggingLogicalErrors)) may provide students with effective debugging strategies to employ when encountering errors. Additionally, features of the programming environment students use could alleviate some of the drivers specific to TBPLs. For example, frame-based environments allow students to write text-based programs while reducing the aforementioned syntax barrier (FrameBasedEditing; StrypePoster2024), and more informative PEMs (e.g., (EffectiveApproachErrorMessageEnhancement; ErrorMessageReadability; HighPrecisionFeedbackSyntaxErrors)) would help reduce the dependency on the teacher for feedback. While the curriculum taught by teachers in our study mandates the teaching of a TBPL, secondary school teachers could instead teach a BBPL as a counter in itself, though this may come at the cost of perceived authenticity (ToBlockOrNotToBlock).
@@
These strategies may take time for students to become accustomed to (CombattingLearnedHelplessness). However, the teachers who report consistently using these counters also report their students becoming more empowered to attempt teacher-independent debugging (CurrentPerspectivesDebugging). Ultimately, this gives the teacher more time to provide more meaningful support to a smaller number of struggling students.
@@
Among the teachers we interviewed, those who reported employing counters tended to be more confident and had decades of experience. This was less true for the less confident teachers, leaving them more susceptible to the consequences of the hands-up problem. More generally, the prevalence of these counters in computing classrooms is unclear. Results from ComputingInCurriculumChallengesStrategies suggest only a fraction of teachers have strategies for teaching debugging and resilience to errors, with a lack of time to teach programming probably a limiting factor.
@@ 5.2 The Implications
Although we only interviewed nine teachers, we expect our findings align with other teachers’ experiences for two reasons. First, previous researchers and teachers have reported struggles in line with the hands-up problem (CombattingLearnedHelplessness; ComputingInCurriculumChallengesStrategies; CurrentPerspectivesDebugging; LackPerserveranceAutonomyDebugging; StumpTheTeacher). Second, the commonality of non-specialist computing teachers and reported lack of content knowledge by some (ExpandingCSChallengesExperiences; ComputingInCurriculumChallengesStrategies) hints at a lack of debugging PCK necessary for effectively supporting students with debugging.
For teachers who report counters to the hands-up problem, experience or professional development seemed to be the reason. This begs the question: how else can counters to the hands-up problem be developed? Or, in other words, how can less experienced or confident teachers teach debugging more confidently? We did not identify any themes related to this in our analysis and thus only provide initial recommendations for future research.
@@
One suggestion is through professional development. Some studies have already reported successful debugging-specific PD, which have been participatory in nature (DoINeedToKnowIfTheTeacher; EffectsOfProfessionalDevelopmentK12Debugging; SupportingStudentsScienceInquiry). However, given that some teachers still report findings aligned with the hands-up problem, debugging-specific PD is either not widespread enough or not being effectively translated into practice. If the former is the case, debugging-related PD needs to be developed or expanded. This does not need to solely focus on debugging; more emphasis on debugging could be included in programming PD workshops, given the intertwined nature of these skills.
Similarly, informing teachers about pedagogical approaches for debugging developed by prior work has huge potential. Although long-term evidence for many debugging-specific teaching approaches is lacking (DebuggingInterventionLitReview), the increasing amount of research in this area at least provides teachers with some suggestions. Some teachers would likely adapt these to their classroom practice (EffectsOfProfessionalDevelopmentK12Debugging), which could provide further insight into the efficacy and applications of these approaches. How teachers can be informed about debugging-specific approaches in a way that encourages classroom adoption is an important research-to-practice consideration, but outside the scope of this article.
These suggestions relate to the improvement of teachers’ ‘debugging PCK’. Despite the challenges debugging presents in programming classrooms, we are not aware of a conceptualisation of debugging PCK in the research literature. A debugging PCK would inevitably have some commonalities with conceptualisations of programming PCK (ProgrammingPCK), such as the motivation for teaching it and common student difficulties. It would also include debugging-specific pedagogy, techniques for effectively managing teacher-student interactions, and other aforementioned counters to the hands-up problem. Therefore, even if a teacher does not have time to explicitly teach the debugging process, a more developed debugging PCK would equip them to deal with the hands-up problem. Once the concept of a debugging PCK is more well-developed, it can be spread through professional development.